# Robust Kernel Hypothesis Testing under Data Corruption

### v1

@@ C.2 Proof of Lemma 2
(i) To simplify the notation, define $U_{0}=T({\widetilde{\mathcal{X}}_{n}})$, $U_{i}=T({\widetilde{\mathcal{X}}_{n}}^{\bm{\pi}_{i}})$ for $i\in[B]$, (i.e., the test statistics based on the uncorrupted dataset ${\widetilde{\mathcal{X}}_{n}}$ and its permuted counterparts ${\widetilde{\mathcal{X}}_{n}}^{\bm{\pi}_{i}}$) and denote the $(1-\alpha)$-quantile of $U_{0},U_{1},\ldots,U_{B}$ as $\widetilde{q}$. Equipped with this notation, observe that $|U_{i}-T_{i}|\leq r\Delta_{T}$ for each $i\in[B]_{0}$, which follows by the repeated use of the triangle inequality and by using the definition of global sensitivity $\Delta_{T}$ (see Equation 1). This inequality yields
Since $U_{0},U_{1},\ldots,U_{B}$ are exchangeable under the null, we have $\mathbb{P}_{P_{0}}(T_{0}>q+2r\Delta_{T})\leq\mathbb{P}_{P}(U_{0}>\tilde{q})\leq\alpha$ for any $P_{0}\in\mathcal{P}_{0}$ [48, Lemma 1], which proves the validity result.
(ii) Using the same notation, we first note that the type II error of the DC test is bounded above as follows:
where we use the fact that $|U_{i}-T_{i}|\leq r\Delta_{T}$ for each $i\in[B]_{0}$. Therefore, in order to prove that the DC test is consistent, it suffices to show that the above upper bound for the type II error converges to zero as $n\rightarrow\infty$ under the given conditions. Now denoting $\overline{U}_{i}=U_{i}+4r\Delta_{T}$ for $i\in[B]_{0}$, by definitions of quantiles, we obtain
$=\mathds{1}(U_{0}>\widetilde{q}+4r\Delta_{T}).$
Using this alternative representation of the test in conjunction with [38, Lemma 8], it follows that
C.3 Proof of Lemma 3
See 3
@@ C.3 Proof of Lemma 3
Starting with dpMMD, we need to verify the condition in Lemma 1 that
$\lim_{n\rightarrow\infty}\mathbb{P}_{P,Q}\biggl{(}\,\widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m})+2\zeta\frac{\sqrt{2K}}{n}\varepsilon^{-1}>\leavevmode\nobreak\ \widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m}^{\bm{\pi}})+2\zeta^{\prime}\frac{\sqrt{2K}}{n}\varepsilon^{-1}\biggr{)}=1.$
Indeed, since $1/(\varepsilon n)\to 0$, it suffices to verify that
$\lim_{n\rightarrow\infty}\mathbb{P}_{P,Q}\Bigl{(}\,\widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m})>\leavevmode\nobreak\ \widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m}^{\bm{\pi}})\Bigr{)}=1,$
which was shown in the proof of [38, Theorem 5]. Therefore, dpMMD is consistent.
Next, in order to prove the consistency of dcMMD, it is enough to verify the condition in Lemma 2:
$\lim_{n\rightarrow\infty}\mathbb{P}_{P_{1}}\biggl{(}\,\widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m})\leavevmode\nobreak\ >\leavevmode\nobreak\ \widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m}^{\bm{\pi}})+4r\frac{\sqrt{2K}}{n}\biggr{)}=1.$
This limit holds since $r/n\to 0$ along with the previous limit observation made for dpMMD. This completes the proof of Lemma 3. ∎
C.4 Proof of Theorem 1
See 1
We separately prove: (a) the uniform separation of dpMMD, (b) the uniform separation of dcMMD, and (c) the optimal uniform separation, in order. We emphasise that from Lemmas 1 and 2, both dpMMD and dcMMD control the non-asymptotic type I error rate at level $\alpha$. Hence, in what follows, we focus on their type II error guarantees.
(a) Proof of uniform separation of dpMMD.
Recall that the dpMMD test with level $\alpha$ is defined with privacy parameters $\varepsilon=r^{-1}\max\!\big{\{}\!\log(e/\alpha),\,\log(e/\beta)\big{\}}$ and $\delta=0$, with adjusted level parameter $\alpha_{\mathtt{dp}}\coloneqq e^{-r\varepsilon}\alpha$. As in the proof of Lemma 1, we may use the DP group property to bound the type II error of dpMMD as
[…]
Now, by leveraging [38, Theorem 7] with $n\asymp m$ and $B\geq 6\alpha_{\mathtt{dp}}^{-1}\log(2\beta_{\mathtt{dp}}^{-1})$, we see that the minimum value of $\rho$ that controls the type II error as
$\sup_{(P,Q)\in\mathcal{P}_{\mathrm{MMD}_{k}}(\rho)}\mathbb{P}_{P,Q}\bigl{(}\textrm{dpMMD fails to reject $\mathcal{H}_{0}$}\,\,|\,\,\textrm{uncorrupted data}\bigr{)}\leq\beta_{\mathtt{dp}},$
satisfies
$\rho\leavevmode\nobreak\ \leq\leavevmode\nobreak$ $C_{K}^{\prime}\max\Biggl{\{}\sqrt{\frac{\max\!\big{\{}\!\log(e/\alpha_{\mathtt{dp}}),\,\log(e/\beta_{\mathtt{dp}})\big{\}}}{n}},\,\frac{\max\!\big{\{}\!\log(e/\alpha_{\mathtt{dp}}),\,\log(e/\beta_{\mathtt{dp}})\big{\}}}{n\varepsilon}\Biggr{\}}$
[…]
(b) Proof of uniform separation of dcMMD.
As shown in Equation 4, the type II error of the DC test based on corrupted data ${\mathcal{X}_{n+m}}$ is bounded above by the type II error of the modified DC test based on uncorrupted data ${\widetilde{\mathcal{X}}_{n+m}}$. This modified DC test uses the cutoff value $\widetilde{q}+4r\Delta_{T}$ instead of $q+2r\Delta_{T}$. This slightly inflated cutoff value only affects the constant factor in the minimum uniform separation. We therefore assume that the data are not corrupted throughout the proof and derive the minimum uniform separation for the dcMMD test.
@@ C.4 Proof of Theorem 1
define the event that $E_{1}:=\{\widetilde{q}\leq q_{\infty,1-\alpha/2}\}$. With $B\geq 3\alpha^{-2}\bigl{\{}\log\bigl{(}8/\beta\bigr{)}+\alpha(1-\alpha)\}$, following the proof of [50, Proposition 4] shows that $\mathbb{P}_{P,Q}(E_{1})\geq 1-\beta/2$. Moreover, [37, Theorem 5.1] with the condition $n\asymp m$ ensures that
@@
$E_{2}:=\biggl{\{}\bigg{|}\mathrm{MMD}_{k}(P,Q)-\widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m})\bigg{|}\leq C_{2,K}\sqrt{\frac{\log(4/\beta)}{n}}\biggr{\}}.$
which satisfies $\mathbb{P}_{P,Q}(E_{2})\geq 1-\beta/2$ by [29, Theorem 7]. With the two events $E_{1}$ and $E_{2}$, holding with high probability, the type II error of the dcMMD test is bounded above as
@@
where the last equality holds by taking $C_{K}>C_{3,K}$ in the theorem statement. ∎
(c) Proof of optimal uniform separation.
We aim to prove that if the separation parameter $\rho$ in Equation 5 is smaller than the following threshold:
@@
no test can have power greater than $1-\beta$ uniformly over $\mathcal{P}_{\mathrm{MMD}_{k}}(\rho)$ where $C_{\eta}>0$ is a small constant depending on $\eta$ in 1. The first part of the lower bound, involving $n^{-1/2}$, was obtained in [38, Appendix E.10.1], except that we have $\log(e/(\alpha+\beta))$ instead of $\log(1/(\alpha+\beta))$. Under the condition $\alpha+\beta<0.4$, it can be seen that both are the same rate by adjusting the constant factor $C_{\eta}$. Hence, it is enough to prove that the second part of the threshold, involving $r/n$.
As remarked in [38, Appendix E.10.1], the minimax separation for two-sample testing is not smaller than that for one-sample testing. Therefore, it suffices to derive the lower bound result for one-sample testing. Specifically, given i.i.d. observations $Y_{1},\ldots,Y_{n}$ drawn from $P$, we are interested in distinguishing $\mathcal{H}_{0}:P=Q_{0}$ against $\mathcal{H}_{1}:\mathrm{MMD}_{k}(P,Q_{0})\geq\rho$ where $Q_{0}$ is a hypothesised (known) distribution, which will be specified later on.
@@
Choose some specific distribution $P_{0}\in\mathcal{P}_{1}(\rho)$ and $M_{r}^{\ast}\in\mathcal{M}_{r}$ specified later on. By Le Cam’s two point method [39, 40], the minimax type II error satisfies
@@
which is guaranteed to exist (minimiser of optimal transport). Here $\pi(P_{0},Q_{0})$ denotes the set of all couplings between $P_{0}$ and $Q_{0}$. Now, following the proof of [3, Theorem 1], take $M_{r}^{\ast}$ as
@@
which corrupts at most $r$ samples by construction, and hence belongs to $\mathcal{M}_{r}$. Hence
@@
where step (i) follows by the coupling lemma of the TV distance, i.e., $d_{\mathrm{TV}}(X,Y)\leq\mathbb{P}(X\neq Y)$, step (ii) uses Markov’s inequality, and step (iii) follows by [2, Lemma 20] along with the definition of $F$ in (6). Therefore, if
@@
then the minimax type II error is bounded below by $\beta$. Finally, as in [38, Appendix E.10.1], we choose $P_{0}=p_{0}\delta_{x}+(1-p_{0})\delta_{v}$ and $Q_{0}=q_{0}\delta_{x}+(1-q_{0})\delta_{v}$, where $x,v\in\mathbb{R}^{d}$, $10<p_{0},q_{0}<$ 1 and $\delta_{x}$ is a Dirac measure at $x$, which yields
@@
Hence the second part of the lower bound holds. This together with the condition $\alpha\asymp\beta$ completes the proof of Theorem 1(ii). ∎
C.5 Proof of Lemma 4
See 4
@@ C.5 Proof of Lemma 4
The proof of Lemma 4 follows a similar approach to that of Lemma 3. Starting with dpHSIC, we need to verify the condition in Lemma 1 that
$\lim_{n\rightarrow\infty}\mathbb{P}_{P_{Y\!Z}}\biggl{(}\,\widehat{\mathrm{HSIC}}(\widetilde{\mathcal{X}}_{n})+2\zeta\frac{4(n-1)\sqrt{KL}}{n^{2}}\varepsilon^{-1}>\leavevmode\nobreak\ \widehat{\mathrm{HSIC}}(\widetilde{\mathcal{X}}_{n}^{\bm{\pi}})+2\zeta^{\prime}\frac{4(n-1)\sqrt{KL}}{n^{2}}\varepsilon^{-1}\biggr{)}=1.$
Indeed, since $1/(\varepsilon n)\to 0$, it suffices to verify that
$\lim_{n\rightarrow\infty}\mathbb{P}_{P_{Y\!Z}}\Bigl{(}\,\widehat{\mathrm{HSIC}}(\widetilde{\mathcal{X}}_{n})>\leavevmode\nobreak\ \widehat{\mathrm{HSIC}}(\widetilde{\mathcal{X}}_{n}^{\bm{\pi}})\Bigr{)}=1,$
which was shown in the proof of [38, Theorem 6]. Therefore, dpHSIC is consistent.
Next, in order to prove the consistency of dcHSIC, it is enough to verify the condition in Lemma 2:
$\lim_{n\rightarrow\infty}\mathbb{P}_{P_{Y\!Z}}\biggl{(}\,\widehat{\mathrm{HSIC}}(\widetilde{\mathcal{X}}_{n})\leavevmode\nobreak\ >\leavevmode\nobreak\ \widehat{\mathrm{HSIC}}(\widetilde{\mathcal{X}}_{n}^{\bm{\pi}})+\frac{16r(n-1)\sqrt{KL}}{n^{2}}\biggr{)}=1.$
This limit holds since $r/n\to 0$ along with the previous limit observation made for dpHSIC. This completes the proof of Lemma 4. ∎
C.6 Proof of Theorem 2
@@
We separately prove: (a) the uniform separation of dpHSIC, (b) the uniform separation of dcHSIC, and (c) the optimal uniform separation, in order. The proof idea is essentially the same as the one used in the proof of Theorem 1. As for the robust MMD tests, we emphasise that from Lemmas 1 and 2, both dpHSIC and dcHSIC control the non-asymptotic type I error rate at level $\alpha$. Hence, in what follows, we focus on their type II error guarantees.
(a) Proof of uniform separation of dpHSIC.
Recall that the dpHSIC test with level $\alpha$ is defined with privacy parameters $\varepsilon=r^{-1}\max\!\big{\{}\!\log(e/\alpha),\,\log(e/\beta)\big{\}}$ and $\delta=0$, with adjusted level parameter $\alpha_{\mathtt{dp}}\coloneqq e^{-r\varepsilon}\alpha$. As in the proof of Lemma 1, we may use the DP group property to bound the type II error of dpHSIC as
[…]
Now, by leveraging [38, Theorem 12] and $B\geq 6\alpha_{\mathtt{dp}}^{-1}\log(2\beta_{\mathtt{dp}}^{-1})$, we see that the minimum value of $\rho$ that controls the type II error as
satisfies
$\rho\leavevmode\nobreak\ \leq\leavevmode\nobreak$ $C_{K,L}^{\prime}\max\Biggl{\{}\sqrt{\frac{\max\!\big{\{}\!\log(e/\alpha_{\mathtt{dp}}),\,\log(e/\beta_{\mathtt{dp}})\big{\}}}{n}},\,\frac{\max\!\big{\{}\!\log(e/\alpha_{\mathtt{dp}}),\,\log(e/\beta_{\mathtt{dp}})\big{\}}}{n\varepsilon}\Biggr{\}}$
[…]
where $C_{K,L}^{\prime},C_{K,L}$ are constants depending on $K$ and $L$. where the last equality holds since we set $\varepsilon=r^{-1}\max\!\big{\{}\!\log(e/\alpha),\,\log(e/\beta)\big{\}}$. This proves the uniform separation of the dpHSIC test.
(b) Proof of uniform separation of dcHSIC.
The proof for the uniform separation of dcHSIC closely resembles that of dcMMD in Theorem 1. The only difference is that we need to use concentration inequalities for the empirical HSIC rather than the empirical MMD. To present details, we assume that the data are not corrupted throughout the proof and derive the minimum uniform separation for dcHSIC. As explained in the proof of Theorem 1, this assumption does not affect the separation rate due to the inequality given in Equation 4.
@@
and define the event that $E_{1}:=\{\widetilde{q}\leq q_{\infty,1-\alpha/2}\}$. With $B\geq 3\alpha^{-2}\bigl{\{}\log\bigl{(}8/\beta\bigr{)}+\alpha(1-\alpha)\}$, we have $\mathbb{P}_{P_{Y\!Z}}(E_{1})\geq 1-\beta/2$, which is shown in the proof of [50, Proposition 4]. Moreover, [38, Lemma 12] ensures that
@@
which holds $\mathbb{P}_{P_{Y\!Z}}(E_{2})\geq 1-\beta/2$ by [38, Lemma 14]. With the two events $E_{1}$ and $E_{2}$, holding with high probability, the type II error of dcHSIC is bounded above as
@@
where the last equality holds by taking $C_{K,L}>C_{4,K,L}$ in the theorem statement. This proves the uniform separation of dcHSIC. ∎
(c) Proof of optimal uniform separation.
We would like to show that if the separation parameter $\rho$ in Equation 8 is smaller than the following threshold:
@@
no test can have power greater than $1-\beta$ uniformly over $\mathcal{P}_{\mathrm{HSIC}_{k,\ell}}(\rho)$ where $C_{\eta_{Y},\eta_{Z}}>0$ is a small constant depending on $\eta_{Y},\eta_{Z}$ in 2. The first part of the lower bound, involving $n^{-1/2}$, was obtained in [38, Appendix F.4.1] with a slight modification to a constant factor $C_{\eta_{Y},\eta_{Z}}$ to replace $\log(1/(\alpha+\beta))$ with $\log(e/(\alpha+\beta))$ where $\alpha+\beta<0.4$. Hence, it is enough to prove the second part of the threshold, involving $r/n$. The proof of this claim is essentially the same as the proof of Theorem 1(ii). To prove the desired result, it hence suffices to find a distribution $P_{Y\!Z}$ such that
@@
To this end, as in [38, Appendix F.4.1], construct the distribution of $(Y,Z)$ as
@@
for some distinct $y_{1},y_{2}\in\mathbb{R}^{d_{Y}}$ and $z_{1},z_{2}\in\mathbb{R}^{d_{Z}}$ such that $y_{1}-y_{2}=y_{0}$ and $z_{1}-z_{2}=z_{0}$, and $\mu\in(0,1/4]$. Following the calculations in [38, Appendix F.4.1], the HSIC of such $P_{Y\!Z}$ and the TV distance between $P_{Y\!Z}$ and $P_{Y}\times P_{Z}$ can be computed as

### v3

@@ C.2 Proof of Lemma 2
In order to prove the consistency of dcMMD, it is enough to verify the condition in Lemma 1:
$\lim_{n\rightarrow\infty}\mathbb{P}_{P_{1}}\biggl{(}\,\widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m})~{}>~{}\widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m}^{\bm{\pi}})+4r\frac{\sqrt{2K}}{n}\biggr{)}=1.$
As $r/n\to 0$, we need to verify that
$\lim_{n\rightarrow\infty}\mathbb{P}_{P,Q}\Bigl{(}\,\widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m})>~{}\widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m}^{\bm{\pi}})\Bigr{)}=1,$
which holds by the result of Kim and Schrab, (2023, Theorem 5). This proves that dcMMD is consistent. ∎
C.3 Proof of Theorem 1
See 1
@@ C.3 Proof of Theorem 1
We emphasise that from Lemma 1, dcMMD controls the non-asymptotic type I error rate at level $\alpha$. Hence, we focus on the type II error guarantees.
(i) As shown in Equation 3, the type II error of the DC test based on corrupted data ${\mathcal{X}_{n+m}}$ is bounded above by the type II error of the modified DC test based on uncorrupted data ${\widetilde{\mathcal{X}}_{n+m}}$. This modified DC test uses the cutoff value $\widetilde{q}+4r\Delta_{T}$ instead of $q+2r\Delta_{T}$. This slightly inflated cutoff value only affects the constant factor in the minimum uniform separation. We therefore assume that the data are not corrupted throughout the proof and derive the minimum uniform separation for the dcMMD test.
@@
define the event that $E_{1}:=\{\widetilde{q}\leq q_{\infty,1-\alpha/2}\}$. With $B\geq 3\alpha^{-2}\bigl{\{}\log\bigl{(}8/\beta\bigr{)}+\alpha(1-\alpha)\}$, following the proof of Schrab et al., (2023, Proposition 4) shows that $\mathbb{P}_{P,Q}(E_{1})\geq 1-\beta/2$. Moreover, Kim, (2021, Theorem 5.1) with the condition $n\asymp m$ ensures that
@@
$E_{2}:=\biggl{\{}\bigg{|}\mathrm{MMD}_{k}(P,Q)-\widehat{\mathrm{MMD}}(\widetilde{\mathcal{X}}_{n+m})\bigg{|}\leq C_{2,K}\sqrt{\frac{\log(4/\beta)}{n}}\biggr{\}},$
which satisfies $\mathbb{P}_{P,Q}(E_{2})\geq 1-\beta/2$ by Gretton et al., 2012a (, Theorem 7). With the two events $E_{1}$ and $E_{2}$, holding with high probability, the type II error of the dcMMD test is bounded above as
@@
where the last equality holds by taking $C_{K}>C_{3,K}$ in the theorem statement.
(ii) We aim to prove that if the separation parameter $\rho$ in Equation 7 is smaller than the following threshold:
@@
no test can have power greater than $1-\beta$ uniformly over $\mathcal{P}_{\mathrm{MMD}_{k}}(\rho)$ where $C_{\eta}>0$ is a small constant depending on $\eta$ in 1. The first part of the lower bound, involving $n^{-1/2}$, was obtained in Kim and Schrab, (2023, Appendix E.10.1), except that we have $\log(e/(\alpha+\beta))$ instead of $\log(1/(\alpha+\beta))$. Under the condition $\alpha+\beta<0.4$, it can be seen that both are the same rate by adjusting the constant factor $C_{\eta}$. Hence, it is enough to prove that the second part of the threshold, involving $r/n$.
As remarked in Kim and Schrab, (2023, Appendix E.10.1), the minimax separation for two-sample testing is not smaller than that for one-sample testing. Therefore, it suffices to derive the lower bound result for one-sample testing. Specifically, given i.i.d. observations $Y_{1},\ldots,Y_{n}$ drawn from $P$, we are interested in distinguishing $\mathcal{H}_{0}:P=Q_{0}$ against $\mathcal{H}_{1}:\mathrm{MMD}_{k}(P,Q_{0})\geq\rho$ where $Q_{0}$ is a hypothesised (known) distribution, which will be specified later on.
@@
Choose some specific distribution $P_{0}\in\mathcal{P}_{1}(\rho)$ and $M_{r}^{\ast}\in\mathcal{M}_{r}$ specified later on. By Le Cam’s two point method (Le Cam,, 1973, 2012), the minimax type II error satisfies
@@
which is guaranteed to exist (minimiser of optimal transport). Here $\pi(P_{0},Q_{0})$ denotes the set of all couplings between $P_{0}$ and $Q_{0}$. Now, following the proof of Acharya et al., 2021b (, Theorem 1), take $M_{r}^{\ast}$ as
@@
which corrupts at most $r$ samples by construction, and hence belongs to $\mathcal{M}_{r}$. Hence, we get
@@
where step (i) follows by the coupling lemma of the TV distance, i.e., $d_{\mathrm{TV}}(X,Y)\leq\mathbb{P}(X\neq Y)$, step (ii) uses Markov’s inequality, and step (iii) follows by Acharya et al., 2021a (, Lemma 20) along with the definition of $F$ in (4). Therefore, if
@@
then the minimax type II error is bounded below by $\beta$. Finally, as in Kim and Schrab, (2023, Appendix E.10.1), we choose $P_{0}=p_{0}\delta_{x}+(1-p_{0})\delta_{v}$ and $Q_{0}=q_{0}\delta_{x}+(1-q_{0})\delta_{v}$, where $x,v\in\mathbb{R}^{d}$, $10<p_{0},q_{0}<$ 1 and $\delta_{x}$ is a Dirac measure at $x$, which yields
@@
Hence, the second part of the lower bound holds. This together with the condition $\alpha\asymp\beta$ completes the proof of Theorem 1(ii). ∎
C.4 Proof of Lemma 3
See 3
@@ C.4 Proof of Lemma 3
The proof of Lemma 3 mirrors the one of Lemma 2. In order to prove the consistency of dcHSIC, it is enough to verify the condition in Lemma 1:
$\lim_{n\rightarrow\infty}\mathbb{P}_{P_{Y\!Z}}\biggl{(}\,\widehat{\mathrm{HSIC}}(\widetilde{\mathcal{X}}_{n})~{}>~{}\widehat{\mathrm{HSIC}}(\widetilde{\mathcal{X}}_{n}^{\bm{\pi}})+\frac{16r(n-1)\sqrt{KL}}{n^{2}}\biggr{)}=1.$
Since $r/n\to 0$, it is enough to verify that
$\lim_{n\rightarrow\infty}\mathbb{P}_{P_{Y\!Z}}\Bigl{(}\,\widehat{\mathrm{HSIC}}(\widetilde{\mathcal{X}}_{n})>~{}\widehat{\mathrm{HSIC}}(\widetilde{\mathcal{X}}_{n}^{\bm{\pi}})\Bigr{)}=1,$
which was shown in the proof of Kim and Schrab, (2023, Theorem 6). Therefore, dcHSIC is consistent. ∎
C.5 Proof of Theorem 2
@@
Proof.
We stress that from Lemma 1, dpHSIC controls the non-asymptotic type I error rate at level $\alpha$, so we focus on the type II error guarantees.
(i) The proof for the uniform separation of dcHSIC closely resembles that of dcMMD in Theorem 1. The only difference is that we need to use concentration inequalities for the empirical HSIC rather than the empirical MMD. To present details, we assume that the data are not corrupted throughout the proof and derive the minimum uniform separation for dcHSIC. As explained in the proof of Theorem 1, this assumption does not affect the separation rate due to the inequality given in Equation 3.
@@
and define the event that $E_{1}:=\{\widetilde{q}\leq q_{\infty,1-\alpha/2}\}$. With $B\geq 3\alpha^{-2}\bigl{\{}\log\bigl{(}8/\beta\bigr{)}+\alpha(1-\alpha)\}$, we have $\mathbb{P}_{P_{Y\!Z}}(E_{1})\geq 1-\beta/2$, which is shown in the proof of Schrab et al., (2023, Proposition 4). Moreover, Kim and Schrab, (2023, Lemma 12) ensures that
@@
which holds with $\mathbb{P}_{P_{Y\!Z}}(E_{2})\geq 1-\beta/2$ by Kim and Schrab, (2023, Lemma 14). With the two events $E_{1}$ and $E_{2}$, holding with high probability, the type II error of dcHSIC is bounded above as
@@
where the last equality holds by taking $C_{K,L}>C_{4,K,L}$ in the theorem statement. This proves the uniform separation of dcHSIC.
(ii) We would like to show that if the separation parameter $\rho$ in Equation 8 is smaller than the following threshold:
@@
no test can have power greater than $1-\beta$ uniformly over $\mathcal{P}_{\mathrm{HSIC}_{k,\ell}}(\rho)$ where $C_{\eta_{Y},\eta_{Z}}>0$ is a small constant depending on $\eta_{Y},\eta_{Z}$ in 2. The first part of the lower bound, involving $n^{-1/2}$, was obtained in Kim and Schrab, (2023, Appendix F.4.1) with a slight modification to a constant factor $C_{\eta_{Y},\eta_{Z}}$ to replace $\log(1/(\alpha+\beta))$ with $\log(e/(\alpha+\beta))$ where $\alpha+\beta<0.4$. Hence, it is enough to prove the second part of the threshold, involving $r/n$. The proof of this claim is essentially the same as the proof of Theorem 1(ii). To prove the desired result, it hence suffices to find a distribution $P_{Y\!Z}$ such that
@@
To this end, as in Kim and Schrab, (2023, Appendix F.4.1), construct the distribution of $(Y,Z)$ as
@@
for some distinct $y_{1},y_{2}\in\mathbb{R}^{d_{Y}}$ and $z_{1},z_{2}\in\mathbb{R}^{d_{Z}}$ such that $y_{1}-y_{2}=y_{0}$ and $z_{1}-z_{2}=z_{0}$, and $\mu\in(0,1/4]$. Following the calculations in Kim and Schrab, (2023, Appendix F.4.1), the HSIC of such $P_{Y\!Z}$ and the TV distance between $P_{Y\!Z}$ and $P_{Y}\times P_{Z}$ can be computed as